OpenStudy (anonymous):

Use a double angle formula the expression a)sinxcosx b)cos^2(x)- 1/2 c) 10sin^2(x)-5 d) (sinx-cosx)(sinx-cosx)

OpenStudy (anonymous):

*rewrite the expression

OpenStudy (cwrw238):

a) use sin2x = 2sinxcosx

OpenStudy (cwrw238):

b) use cos2x = 2cos^2x - 1

OpenStudy (jdoe0001):

recall that -> sin(2x)=2sin(x)cos(x) so what you'd do is, make the right-hand-side look like sin(x)cos(x) by simplifying so what do you think you could use to make \(\bf {\color{blue}{ sin(2x)}}=2sin(x)cos(x)\implies {\color{blue}{ \square ?}}=sin(x)cos(x)\quad ?\)

OpenStudy (anonymous):

2

OpenStudy (jdoe0001):

times 2? over 2? under 2? + 2?

OpenStudy (anonymous):

times 2...maybe?

OpenStudy (jdoe0001):

let's see that \(\bf {\color{blue}{ sin(2x)}}=2sin(x)cos(x)\implies 2{\color{blue}{ sin(2x)}}=2\cdot 2sin(x)cos(x) \\ \quad \\ \implies 2{\color{blue}{ sin(2x)}}=4sin(x)cos(x)\) nope, the left-side doesn't look like sin(x)cos(x)

OpenStudy (jdoe0001):

right-hand-side rather

OpenStudy (anonymous):

so times 1?

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin(2x)}}=2sin(x)cos(x)\implies 1{\color{blue}{ sin(2x)}}=1\cdot 2sin(x)cos(x) \\ \quad \\ \implies 2{\color{blue}{ sin(2x)}}=2sin(x)cos(x)\) no dice on that either, the right-side doesn't look like sin(x)cos(x) yet

OpenStudy (anonymous):

okay, I give up

OpenStudy (jdoe0001):

woops, darn typos... anyhow \(\bf {\color{blue}{ sin(2x)}}=2sin(x)cos(x)\implies 1{\color{blue}{ sin(2x)}}=1\cdot 2sin(x)cos(x) \\ \quad \\ \implies {\color{blue}{ sin(2x)}}=2sin(x)cos(x)\) still though, no dice on that one

OpenStudy (jdoe0001):

lemme rewrite it a bit differently

OpenStudy (jdoe0001):

actually... lemme show it instead... let use DIVIDE by 2 both sides \(\bf {\color{blue}{ sin(2x)}}=2sin(x)cos(x)\implies \cfrac{{\color{blue}{ sin(2x)}}}{2}=\cfrac{\cancel{2}sin(x)cos(x)}{\cancel{2}}\) what do you think?

OpenStudy (anonymous):

i'm following

OpenStudy (jdoe0001):

well, what would the right-side give you?

OpenStudy (anonymous):

sin(x)cos(x) right?

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin(2x)}}=2sin(x)cos(x)\implies \cfrac{{\color{blue}{ sin(2x)}}}{2}=\cfrac{\cancel{2}sin(x)cos(x)}{\cancel{2}}\\ \quad \\ \implies \cfrac{{\color{blue}{ sin(2x)}}}{2}=sin(x)cos(x)\)

OpenStudy (anonymous):

okay so what about b?

OpenStudy (jdoe0001):

the b) one, as cwrw238 already pointed out \(\bf {\color{blue}{ cos(2x)}}=2cos^2(x)-1\implies {\color{blue}{ \square ?}}=cos^2(x)-\frac{1}{2}\) so you'd simplify till the right-hand-side looks like that :) try also dividing both sides by 2, that is \(\bf \cfrac{{\color{blue}{ cos(2x)}}}{2}=\cfrac{2cos^2(x)-1}{2}\)

OpenStudy (anonymous):

You're a life-saver thank you

OpenStudy (anonymous):

wait, what about d?